# MUBen: Benchmarking the Uncertainty of Molecular Representation Models

### v3

@@ Appendix E Additional Results
In this section, we supplement the discussions from § 5 with further results and analyses. Given the extensive volume of outcomes from our experiments, we have consolidated them into tables at the end of this article. For enhanced accessibility and offline analysis, we have also made these results available as CSV files in our GitHub repository.
@@ E.1 More Visualization of Uncertainty Estimation Performance
E.2 Complete Fine-Tuning Results
@@
Challenges are encountered during training on the notably imbalanced HIV and MUV dataset, depicted in Table 4. SGLD at times proves unsuccessful due to weak directional cues from the true labels, overwhelmed by the random noise embedded in the Langevin dynamics. With TorchMD-NET backbone, SGLD even fails training and constantly produces NaN values regardless of the hyper-parameters we choose. Conversely, the application of Focal Loss demonstrates effective performance on the MUV dataset, aligning with assertions in Lin et al. (2017). ECE also may fail to compute on these datasets as predicted probabilities $\hat{p}$ are excessively minimal, leading to unpopulated bins and subsequent “division by zero” errors.
In terms of the majority of datasets, the performance trends closely adhere to our previous discussions. However, an anomaly is observed with the BBBP and ClinTox datasets, where ChemBERTa produces higher performance than the more powerful models, deviating from prior findings (Ahmad et al., 2022; Zhou et al., 2023). Conceptually, the ChemBERTa checkpoint we use is solely pre-trained on the MLM objective, which is unrelated to downstream tasks and should not include label information. Consequently, we hypothesize that either the SMILES strings are occasionally more expressive than other descriptors with respect to the physiological properties, or the checkpoint may have undergone fine-tuning prior to upload, resulting in potential contamination.
Table 7: Averaged ranks of UQ methods across all datasets and backbone models. Lower value indicates better resuls. Temperature Scaling and Focal Loss are not applicable to regression tasks.
@@ E.3 Results on Random Split Datasets
Tables 23 to 30 illustrate the efficacy of primary backbone models across datasets with random divisions. Mirroring the discussion in § E.1, Table 8 and Figure 7 depict the average (reciprocal) rankings of UQ methods. For the most part, the comparative performance among the backbone models and UQ methods is consistent with those observed in scaffold split datasets. Nonetheless, certain exceptions emerge, especially concerning BBP, SGLD, and SWAG. On randomly split classification datasets, both BBP and SGLD continue to underperform, while a marked decline is noted in their regression NLL. This phenomenon might be attributed to the balanced nature of random split datasets, intensifying the predictive indications from actual labels. Given this environment, the noise from Langevin dynamics or Gaussian samples, which usually aids generalization, might become detrimental. This revelation highlights the capability of traditional Bayesian methods in aptly navigating OOD data (i.e., addressing epistemic uncertainty), mainly through curbing overfitting. However, they may struggle when dealing with in-distribution attributes (i.e., addressing aleatoric uncertainty) having relatively stable labels in the context of molecular property prediction. On the other hand, SWAG, another approximation of the Bayesian network that is almost useless on OOD data, delivers a much more impressive performance on random split datasets. Deep Ensembles, MC Dropout, and Temperature Scaling still emerge as the most reliable UQ methods, although Deep Ensembles no longer dominates the benchmark in the in-distribution case.
@@
Table 8: Averaged ranks of UQ methods across all randomly split datasets.
@@ E.4 Results with Frozen Backbone Models
Analysis of Tables 31 through 38 reveals that when the network parameters of the backbone models are fixed, all models underperform significantly. Among the three primary backbone models, Uni-Mol experiences the most drastic decline, emerging as the least effective in the majority of evaluations. These findings highlight that these backbone models are not inherently optimized for mere feature extraction. The fine-tuning process is pivotal for ensuring optimal model performance, irrespective of the use of UQ methods.

### v4

@@ Appendix E Additional Results
In this section, we supplement the discussions from § 5 with further results and analyses. Given the extensive volume of outcomes from our experiments, we have consolidated them into tables at the end of this article. For enhanced accessibility and offline analysis, we have also made these results available as CSV files in our GitHub repository. [^10] https://github.com/Yinghao-Li/MUBen/tree/main/reports.
@@ E.1 More Visualization of Uncertainty Estimation Performance
E.2 Comprehensive Fine-Tuning Results
@@
Challenges are encountered during training on the notably imbalanced HIV and MUV dataset, depicted in Table 4. SGLD at times proves unsuccessful due to weak directional cues from the true labels, overwhelmed by the random noise embedded in the Langevin dynamics. With TorchMD-NET backbone, SGLD even fails training and constantly produces NaN values regardless of the hyperparameters we choose. Conversely, the application of Focal Loss demonstrates effective performance on the MUV dataset, aligning with assertions in Lin et al. (2017). ECE also may fail to compute on these datasets as predicted probabilities $\hat{p}$ are excessively minimal, leading to unpopulated bins and subsequent “division by zero” errors.
In terms of the majority of datasets, the performance trends closely adhere to our previous discussions. However, an anomaly is observed with the BBBP and ClinTox datasets, where ChemBERTa produces higher performance than the more powerful models, deviating from prior findings (Ahmad et al., 2022; Zhou et al., 2023). Conceptually, the ChemBERTa checkpoint we use is solely pre-trained on the MLM objective, which is unrelated to downstream tasks and should not include label information. Consequently, we hypothesize that either the SMILES strings are potentially more expressive than other descriptors with respect to the physiological properties given that the ChemBERTa’s architectures is not considered as more expressive than GROVER or Uni-Mol.
Table 8: Averaged ranks of UQ methods across all scaffold-split datasets and backbone models. Lower value indicates better resuls. Temperature Scaling and Focal Loss are not applicable to regression tasks.
@@ E.3 Results on Random Split Datasets
Tables 24 to 31 illustrate the efficacy of primary backbone models across datasets with random divisions. Mirroring the discussion in § E.1, Table 9 and Figure 8 depict the average (reciprocal) rankings of UQ methods. For the most part, the comparative performance among the backbone models and UQ methods is consistent with those observed in scaffold split datasets. Nonetheless, certain exceptions emerge, especially concerning BBP, SGLD, and SWAG. On randomly split classification datasets, both BBP and SGLD continue to underperform, while a marked decline is noted in their regression NLL. This phenomenon might be attributed to the balanced nature of random split datasets, intensifying the predictive indications from actual labels. Given this environment, the noise from Langevin dynamics or Gaussian samples, which usually aids generalization, might become detrimental. This revelation highlights the capability of traditional Bayesian methods in aptly navigating OOD data (i.e., addressing epistemic uncertainty), mainly through curbing overfitting. However, they may struggle when dealing with in-distribution attributes (i.e., addressing aleatoric uncertainty) having relatively stable labels in the context of molecular property prediction. On the other hand, SWAG, another approximation of the Bayesian network that is almost useless on OOD data, delivers a much more impressive performance on random split datasets. Deep Ensembles, MC Dropout, and Temperature Scaling still emerge as the most reliable UQ methods, although Deep Ensembles no longer dominates the benchmark in the in-distribution case.
@@
Table 9: Averaged ranks of UQ methods across all randomly split datasets. Lower rank indicates better performance.
@@ E.4 Results with Frozen Backbone Models
Analysis of Tables 32 through 39 reveals that when the network parameters of the backbone models are fixed, all models underperform significantly. Among the three primary backbone models, Uni-Mol experiences the most drastic decline, emerging as the least effective in the majority of evaluations. These findings highlight that these backbone models are not inherently optimized for mere feature extraction. The fine-tuning process is pivotal for ensuring optimal model performance, irrespective of the use of UQ methods.
(a) DNN RMSE $\downarrow$
(b) DNN MAE $\downarrow$
(c) DNN NLL $\downarrow$
(d) DNN CE $\downarrow$
(e) ChemBERTa RMSE $\downarrow$
(f) ChemBERTa MAE $\downarrow$
(g) ChemBERTa NLL $\downarrow$
(h) ChemBERTa CE $\downarrow$
(i) GROVER RMSE $\downarrow$
(j) GROVER MAE $\downarrow$
(k) GROVER NLL $\downarrow$
(l) GROVER CE $\downarrow$
(m) Uni-Mol RMSE $\downarrow$
(n) Uni-Mol MAE $\downarrow$
(o) Uni-Mol NLL $\downarrow$
(p) Uni-Mol CE $\downarrow$
Figure 10: Performance of backbone and UQ methods according to the similarity between test data points and training scaffolds in the QM9 dataset.